# Predictive Pattern Recognition Techniques Towards Spatiotemporal Representation of Plant Growth in Simulated and Controlled Environments: A Comprehensive Review

### v2

@@ 1 Introduction
With the growing prevalence of challenges in global food systems, including restricted access to fresh produce for remote populations and issues faced within traditional cultivation practices, outlooks on controlled environment agriculture (CEA) present potential solutions through precision methods of localized food production [Mirzabaev2023, OSULLIVAN2019133, KurtBenke10.1080/15487733.2017.1394054]. Controlled environments within crop production systems provide opportunities to observe and cultivate plants with diverse characteristics, or phenotypes, making it suitable for specialized applications from trait selection for cultivar development and biopharming [Halperin/10.1111/tpj.13425, HENRY2020168, langstroff2022opportunities, agronomy13030858, YANG2023102267], to food production in extreme and resource scarce environments [NGUYEN2023708, 10.3389/fpls.2020.00801, 10.3389/fpls.2020.00656]. Moreover, the ability to accurately predict and represent the phenotypic expression of a plant and its trajectory at various points in its growth cycle demonstrates a fundamental step towards advancing simulation and modeling research within CEA.
@@
Recent developments in computational and algorithmic theories, coupled with innovations in data collection and fusion techniques, have opened new avenues for exploring plant phenotypes and growth patterns within simulated and controlled environments [10.1111/tpj.14722]. These environments allow for the precise manipulation of variables such as light, temperature, humidity, and nutrient availability, enabling researchers to study the dynamic interactions that influence plant development. By leveraging techniques in predictive pattern recognition, models can be developed to accurately capture and represent plant growth behaviors.
1.1 Background
@@
Predictive models within plant research are often dominated by a frequentist approach to statistical inference, where empirical observations and static conditions are tested against hypotheses to determine the probable outcome through confidence intervals or point estimations. This statistical approach is favored within the field for its simplicity in analysis and interpretation. Frequentist methods, including various forms of regression modeling, provide robust tools for analyzing experimental data and have become the standard in plant research due to their well-established procedures and ease of use [hoshmand2018design, bender2020statistical]. However, this experiment-centric approach is often limited when applied toward plant observation tasks that involve complex data structures, as with analyzing sequential and spatiotemporal images collected within the domain of high-throughput plant phenotyping (HTPP). Moreover, plant growth modeling and forecasting is an ill-posed problem where the trajectory of a plant’s development does not have a unique solution. Dynamic growing conditions and stochastic elements within plant development are prevalent due to the inherent variability and uncertainty that exists in biological systems. For example, plant growth is influenced by numerous interacting factors, including genetic, environmental, and management variables, which can lead to highly complex and non-linear relationships. Thus methods for regularization of the problem should be explored.
@@
Within the context of simulations, computational models aim to predict plant traits, such as biomass accumulation, crop yield, presence of diseases, and overall growth dynamics, during various development stages under varying environmental conditions [bdcc5010002]. Some classical methods have approached this task through a combination of process models that simulate physiological principles and data-driven empirical models, typically obtained through a frequentist framework, to form various subsystems of an overarching hybrid model, offering a detailed understanding of how plants respond to different environmental stimuli at a field scale. Models such as the Crop Environment Resource Synthesis (CERES) [BASSO201627] and the Decision Support System for Agrotechnology Transfer (DSSAT) have been extensively used to forecast plant growth by integrating weather data, soil conditions, and crop management practices [JONES2003235, NGUESSAN2023276]. These models have proven instrumental in evaluating the potential impacts of climate change on crop productivity, guiding decision-making in traditional agriculture. However, these methods often demonstrate the various limitations of a experiment-centric approach.
@@
The first limitation of this classical framework is that the underlying inference models are typically discriminative. The parameters of these models are treated as fixed variables that characterize boundaries used to determine the most probable outcome based on the frequencies of occurrence in sample data. This lack of flexibility in describing data distributions makes these models less effective in representing systems that involve stochastic processes, as with the various biological processes that constitute plant growth. As a result, the deterministic process models derived from these empirical outcomes are often sensitive to initial conditions and are constrained to rigid scopes of inference, making it unsuitable for distant future predictions and the synthesis of realistic data. The second limitation is that the outputs of such models are typically expressed through low-dimensional or scalar values, such as yield, biomass, and vegetative index. While this allows for simple comparative analysis and assessments of seasonality of data points, insights toward spatial information and links between input conditions and plant structural developments are lost. The third limitation is that frequentist-based inference models do not intrinsically incorporate mechanisms to update parameters and integrate new information dynamically. This paradigm makes these models less effective for sequence analysis, as they provide no insights on prior information and treat parameters and datasets as static. In contrast, probabilistic approaches, such as those based on Bayesian inference (BI), explicitly quantify uncertainties and dynamically update with new data [mackay2003information]. The static nature of the classical approach to modeling plant growth can lead to the underestimation of uncertainty, as the underlying assumptions about the ”true” distribution of data may change, compounded by the cumulative effect of decisions made during earlier stages of inference in sequence analysis [mcelreath2018statistical]. An overview of these dynamics between modeling paradigms are illustrated in Figure 1.
@@
In this review, we explore these paradigms and assess the body of works available for various modeling tasks that require spatiotemporal analysis and pattern recognition of plant traits during its growth and development. We further propose insights to advance future works within this field and provide a perspective towards a model-centric probabilistic approach to plant growth pattern recognition and modeling through structured 2D and 3D data representations.
@@ 2 Rationale and Approach
Recent state-of-the-art (SOTA) developments made within the field of HTPP have led to the establishment of data collection platforms and protocols for multimodal datasets, which include image sensor and 3D point cloud data. [walter2015plant, 10.3389/fbioe.2020.623705, saric2022applications, plants11172199, WEN2023165626]. With the introduction of these complex datasets and multi-modal data fusion techniques, plant phenomic research has shifted towards computer vision (CV) and sensor-based methods for efficient assessment of plant traits when compared to traditional, labor-intensive approaches [10.3389/fpls.2019.00508, 10.1093/jxb/erz406]. Advances in CV, particularly through the use of deep learning (DL) models such as convolutional neural networks (CNN), have enabled the extraction of phenotypic traits such as leaf area, plant height, and disease symptoms from high-resolution images with good accuracy [singh2016machine]. 3D imaging techniques, including structure-from-motion (SfM), stereo vision, and light detection and ranging (LiDAR) scanning, provide methods for processing spatial information to produce reconstructed 3D representation of plant architectures [SUN2020195].
While these datasets are applied towards methods for assessing crop quality and quantifying biophysical properties throughout various stages of a plant’s growth cycle, the scope of phenomics research is often limited to classification tasks such as detection and segmentation, with limited work done on CV-based predictive forecasting. In contrast, this review focuses on work that explore the spatiotemporal components of plant phenomics research and the representation of plant traits through time.
@@ 2.1 Literature Overview
A rigorous screening process was applied to articles indexed within the following publication databases; SCOPUS, Web of Science (WOS), PubMed, and Google Scholar. A total of 50 method papers on plant growth modeling and 22 temporal dataset papers were selected and reviewed (Figure 2). These papers include articles published within the last 10 years, 2015-2025, with a particular focus on assessing plant traits over time, where temporal analysis is a key component to the method or dataset. Articles in which traits are observed sparsely over time, for example traits observed once at the beginning of an experiment and once after harvest, are excluded from this review, opting for consistent, relatively high temporal resolution tracking of traits over time.
Figure 2: Breakdown of publications reviewed, categorized by modeling approaches and dataset types. The figure highlights the distribution of works across regression models, latent representation models, functional-structural plant models, and computer vision-based techniques, as well as their focus on temporal scalar, 2D image-based, and 3D point cloud datasets.
2.1.1 Plant Growth Representations and Datasets
@@
2.1.2 Limitations
Due to the exploratory nature of this review, the number of works available for a full systemic review of novel and SOTA techniques within this developing field is limited. This review is constrained to published research and datasets, which may lead to an inherent bias towards studies and datasets that have achieved publication rather than those that may be equally relevant but remain in preprint stages, unpublished, or private.
2.2 Scope of this Review
This review serves as a comprehensive examination of the field of pattern recognition in plant growth. We offer a prospective that utilizes the available SOTA methods towards a probabilistic approach through image-based simulation and generative modeling. The review is organised as such:
  • We define plant growth modeling through various temporal paradigms and examine the prominent methods explored for scalar traits.
  • We examine the application of these modeling paradigms towards 2D and 3D representations.
  • We explore progress made within the developing field of probabilistic and generative computer-vision that utilize image-based forecasting.
  • We provide a prospective towards the current state of research and propose a path for future works within the field.
@@ 2.2 Scope of this Review
Sun et al. (Soybean-MVS) [Sun2023soybeanmvs] Soybean (Glycine max var. DN251, DN252, DN253, HN48, HN51) Segmented plant structure (leaf shape, stem, branches) Point cloud (3D)
Schunck et al. (Pheno4D) [10.1371/journal.pone.0256340] Maize (Zea mays), Tomato (Solanum lycopersicum) Segmented plant structure (leaf shape, stem, branches) Point cloud (3D)
Sultan et al. (Shoots and Roots) [doi:10.34133/2020/2073723] Maize (Zea mays), Tomato (Solanum lycopersicum), rice (Azucena, Bala), Arabidopsis (Arabidopsis thaliana) Root system traits and structure Point cloud (3D)
Uchiyama et al. (KOMATSUNA) [Uchiyama2017Komatsuna] Komatsuna (Brassica rapa var. perviridis) Plant structure Multi-view RGBD images and annotated masks (2D/3D)
Jayasuriya et al. [JAYASURIYA2024110735] Capsicum (Capsicum annuum L.) Plant structure RGBD images (2D/3D)
Kierdorf et al. (GrowliFlower) [Uchiyama2017Komatsuna] Cauliflower (Brassica oleracea var. botrytis) Plant structure RGB and multi-spectral images (2D)
Shadrin et al. [Shadrin2018InstanceSF] Lettuce (Lactuca sativa) Leaf growth and leaf segmentation Top-down RGB images and annotated masks (2D), and environmental sensor readings (tabular)
Genze et al. [genze2024manually] Maize (Zea mays L.), Sorghum (Sorghum bicolor) Plant structure Top-down RGB images and annotated masks (2D)
Thoday-Kennedy et al. [THODAYKENNEDY2023108787] Safflower (Carthamus tinctorius L.) Plant morphology and salt response Multi-view RGB images (2D)
Lärm et al. [larm2023multi] Maize (Zea mays), Wheat (Triticum aestivum) Root system traits and structure RGB images and ground-penetrating radar map (2D), soil sensor data (tabular)
Parmentier et al. [essd-13-3593-2021] General vegetation in Arctic ecosystems Vegetative phenology and climate interactions RGB images (2D) and environmental vegetative index data (tabular)
Craig et al. [craig2022hyperspectral] Maize (Zea mays L.) Various agronomic traits and vegetative index Scalar (tabular)
Dépigny et al. (Plantain-Optim) [DEPIGNY2018671] Plantain (Musa acuminata × Musa balbisiana) Various agronomic traits and biomass Scalar (tabular)
Ngaba et al. [NGABA2023109244] Sugarcane (Saccharum officinarum) Various agronomic traits and biomass Scalar (tabular)
Newman et al. [newman2021multiple] Multiple field-crop species Long-term agronomic traits and biomass Scalar (tabular)
Campbell et al. [CAMPBELL2021107600] Maize (Zea mays) Photosynthetic function and vegetative index Scalar (tabular)
Manjarrez-Sanchez et al. [MANJARREZSANCHEZ2020106274] Greenhouse Tomato (Lycopersicum esculentum) Photosynthetic function and agronomic traits Scalar (tabular)
Niu et al. [essd-14-2851-2022] Maize (Zea mays) Long-term phenological and climate interactions Scalar (tabular)
Zhu et al. [essd-16-277-2024] Multiple field-crop species Long-term phenological and climate interactions Scalar (tabular)
Mergner et al. [mergner2020proteomic] Arabidopsis (Arabidopsis thaliana) Gene expression Scalar (tabular)
Castro et al. [CASTRO2020105834] Camu-camu (Myrciaria dubia) Gene expression Scalar (tabular)
Peng et al. [peng2024transcriptome] Peanut (Arachis hypogaea) Gene expression Scalar (tabular)

### v3

# Predictive Modeling, Pattern Recognition, and Spatiotemporal Representations of Plant Growth in Simulated and Controlled Environments: A Comprehensive Review

@@ 1 Introduction
With the growing prevalence of challenges in global food systems, including restricted access to fresh produce in remote areas and issues faced within traditional cultivation practices, outlooks on controlled environment agriculture (CEA) offer prospective solutions through precision methods of localized food production [1, 2, 3]. Controlled environments within crop production systems provide opportunities to observe and cultivate plants with diverse characteristics, or phenotypes, making it suitable for specialized applications from trait selection for cultivar development and biopharming [4, 5, 6, 7, 8], to food production in extreme and resource scarce environments [9, 10, 11]. Subsequently, the ability to accurately predict and represent the phenotypic expression of a plant and its trajectory at various points in its growth cycle demonstrates a fundamental step towards advancing simulation and modeling research within CEA.
@@
Recent developments in computational and algorithmic theories, coupled with innovations in data collection and fusion techniques, have opened new avenues for exploring plant phenotypes and growth patterns within simulated and controlled environments [12]. These environments allow for the precise manipulation of variables such as light, temperature, humidity, and nutrient availability, enabling researchers to study the dynamic interactions that influence plant development. By leveraging techniques in predictive pattern recognition, models can be developed to accurately capture and represent plant growth behaviors.
1.1 Comparing Controlled and Open Field Environments
To contextualize the focus of this review, it is necessary to compare the differences in observational parameters between controlled environment and open field systems. Open field environments, as with traditional agricultural cropping systems, are characterized by the exposure to highly variable and often uncontrollable factors including precipitation, solar conditions, climate fluctuations, and biotic stressors such as pest, pathogens, and soil microbe root interactions. These fluctuations introduce noise and confounding variables into experimental data, making it difficult to isolate causal relationships in plant development. Moreover, instrumentation methods for capturing these factors of variation (FOV) are often developed for constrained experiments, where samples are collected manually, and remain difficult to scale for application in traditional agricultural systems. Thus, outlooks within open-field research often highlight the bottlenecks in sensing and data handling which in turn limit the assessment of physiological development and crop forecasting to sparse spatiotemporal samples of FOV [13, 14]. However, due to the desire for open environment resilient plant varieties for outdoor cropping systems, trends in field-based phenomics research have primarily focused on phenotype-genotype mapping for applications in breeding programs within open-field environments [15, 16, 17].
Controlled environments are typically found within CEA systems such as greenhouses, vertical farms, and growth chambers and are characterized by a confined growing space in which the environmental conditions are known, regulated, or fully controlled. These controlled environment conditions give rise to regularized observations made on plant behavior by constraining some FOVs within known parameterized boundaries. While not all environmental conditions can be fully isolated, key factors such as light (intensity, duration, spectrum), air and root-zone temperature, humidity, CO2 levels, and nutrient delivery are routinely monitored and controlled [18]. Controlled environments lend themselves well to modeling systems as instrumentation and control systems are easily integrated and necessary to maintain optimal conditions for plant growth. These measurements can then be parameterized to develop advanced representation models for predictive forecasting and simulation.
1.2 Defining Simulated Environments and Plant Growth
The phenotypic outcomes of a plant are influenced by high-dimensional interactions between genotype, environment, and management (G×$\times$×E×$\times$×M), often resulting in complex non-linear responses [19, 20, 21]. Subsequently, the objective of controlled environment systems is to constrain and reduce the influence of the environmental dimension by maintaining known optimal conditions of various crop varieties. However, a second useful objective of these systems becomes apparent as advances in automated sensing and controls are implemented within CEA; The development of robust simulations and plant growth prediction models.
The term simulation can take on two meanings within this context. The first, is an empirical simulation where CEA is applied to mimic open-field conditions such that, the plant’s behavior can be explored under controlled conditions, and the insights to be employed under real world field conditions. The second, is a computational simulation, or in silico representation and quantification of plant behaviors given simulated conditions through models. In this review, the use of the term simulation borrows from both meanings and describes in silico models that predict plant growth under real world conditions, which are developed by observing plant behaviors under controlled environments.
With this definition, it is also important to define plant growth within these models. Plant growth refers to the manifestation of plant traits, or phenotypes, over time. More specifically, the size, mass, and complexity of a plant over time, resulting from cell division, cell expansion, and differentiation processes that are regulated by internal genetic programs and modulated by external environmental factors [22, 23]. Quantitatively, plant growth can be represented by morphological traits such as leaf area, height, and biomass, physiological measurements such as photosynthetic rate and transpiration, or biochemical markers, all of which vary across developmental stages and are sensitive to both spatial and temporal heterogeneity [24]. Moreover, the outputs of these simulations can have various representations from low dimensional quantitative traits to high dimensional 2D and 3D renderings and are explored further in this review.
1.3 Background: Predictive Plant Growth Modeling
@@
Predictive models within plant research are often dominated by a frequentist approach to statistical inference, where empirical observations and static conditions are tested against hypotheses to determine the probable outcome through confidence intervals or point estimations. This statistical approach is favored within the field for its simplicity in analysis and interpretation. Frequentist methods, including various forms of regression modeling, provide robust tools for analyzing experimental data and have become the standard in plant research due to their well-established procedures and ease of use [25, 26]. However, this experiment-centric approach is often limited when applied toward plant observation tasks that involve complex data structures, as with analyzing sequential and spatiotemporal images collected within the domain of high-throughput plant phenotyping (HTPP). Moreover, plant growth modeling and forecasting is an ill-posed problem where the trajectory of a plant’s development does not have a unique solution. Dynamic growing conditions and stochastic elements within plant development are prevalent due to the inherent variability and uncertainty that exists in biological systems. For example, plant growth is influenced by numerous interacting factors, including G×$\times$×E×$\times$×M variables. Thus, methods for regularization of the problem should be explored.
@@
Within the context of simulations, computational models aim to predict plant traits, such as biomass accumulation, crop yield, presence of diseases, and overall growth dynamics, during various development stages under varying environmental conditions [27]. Some classical methods have approached this task through a combination of process models that simulate physiological principles and data-driven empirical models, typically obtained through a frequentist framework, to form various subsystems of an overarching hybrid model, offering a detailed understanding of how plants respond to different environmental stimuli at a field scale. Models such as the Crop Environment Resource Synthesis (CERES) [28] and the Decision Support System for Agrotechnology Transfer (DSSAT) have been extensively used to forecast plant growth by integrating weather data, soil conditions, and crop management practices [29, 30]. These models have proven instrumental in evaluating the potential impacts of climate change on crop productivity, guiding decision-making in traditional agriculture. However, these methods often demonstrate the various limitations of a experiment-centric approach.
@@
The first limitation of this classical framework is that the underlying inference models are typically discriminative. The parameters of these models are treated as fixed variables that characterize boundaries used to determine the most probable outcome based on the frequencies of occurrence in sample data. This lack of flexibility in describing data distributions makes these models less effective in representing systems that involve stochastic processes, as with the various biological processes that constitute plant growth. As a result, the deterministic process models derived from these empirical outcomes are often sensitive to initial conditions and are constrained to rigid scopes of inference, making it unsuitable for distant future predictions and the synthesis of realistic data. The second limitation is that the outputs of such models are typically expressed through low-dimensional or scalar quantitative traits, such as yield, biomass, and vegetative index. While this allows for simple comparative analysis and assessments of seasonality of data points, insights toward spatial information and links between input conditions and plant structural developments are lost. The third limitation is that frequentist-based inference models do not intrinsically incorporate mechanisms to update parameters and integrate new information dynamically. This paradigm makes these models less effective for sequence analysis, as they provide no insights on prior information and treat parameters and datasets as static. In contrast, probabilistic approaches, such as those based on Bayesian inference (BI), explicitly quantify uncertainties and dynamically update with new data [31]. The static nature of the classical approach to modeling plant growth can lead to the underestimation of uncertainty, as the underlying assumptions about the ”true” distribution of data may change, compounded by the cumulative effect of decisions made during earlier stages of inference in sequence analysis [32]. An overview of these dynamics between modeling paradigms are illustrated in Figure 1.
1.4 Scope of this Review
@@
We explore these paradigms and assess the body of works available for various modeling tasks that require spatiotemporal analysis and pattern recognition of plant traits during its growth and development. We further propose insights to advance future works within this field and provide a perspective towards a model-centric probabilistic approach to plant growth pattern recognition and modeling through structured 2D and 3D data representations. This review serves as a comprehensive examination of the field of pattern recognition in plant growth. The review is organized as follows:
  • We present our rational and approach for presenting the works and state our limitations. (Section 2)
  • We define plant growth modeling through various temporal paradigms and examine the prominent methods explored for scalar quantitative traits. (Section 3)
  • We examine the application of these modeling paradigms towards 2D and 3D representations. (Section 4)
  • We explore progress made within the developing field of probabilistic and generative computer-vision that utilize image-based forecasting. (Section 4.2)
  • We provide a prospective towards the current state of research and propose a path for future works within the field. (Section 5)
@@ 2 Rationale and Approach
Recent state-of-the-art (SOTA) developments made within the field of HTPP have led to the establishment of data collection platforms and protocols for multimodal datasets, which include image sensor and 3D point cloud data [33, 34, 35, 36, 37]. With the introduction of these complex datasets and multi-modal data fusion techniques, plant phenomic research has shifted towards computer vision (CV) and sensor-based methods for efficient assessment of plant traits when compared to traditional, labor-intensive approaches [38, 39]. Advances in CV, particularly through the use of deep learning (DL) models such as convolutional neural networks (CNN), have enabled the extraction of phenotypic traits such as leaf area, plant height, and disease symptoms from high-resolution images with good accuracy [40]. 3D imaging techniques, including structure-from-motion (SfM), stereo vision, and light detection and ranging (LiDAR) scanning, provide methods for processing spatial information to produce reconstructed 3D representation of plant architectures [41].
While these datasets are applied towards methods for assessing crop quality and quantifying biophysical properties throughout various stages of a plant’s growth cycle, the scope of phenomics research is often limited to recognition tasks such as detection and segmentation, with limited work done on CV-based predictive forecasting. In contrast, this review focuses on work that explore the spatiotemporal components of plant phenomics research and the representation of plant traits through time.
@@ 2.1 Literature Overview
A rigorous screening process was applied to articles indexed within the following publication databases: SCOPUS, Web of Science (WOS), PubMed, and Google Scholar. A total of 51 method papers on plant growth modeling and 22 temporal dataset papers were selected and reviewed (Figure 2). These papers include articles published between the years, 2015-2025, with a particular focus on assessing plant traits over time, where temporal analysis is a key component to the method or dataset. Foundational works predating 2015, such as standard numerical crop models, are referenced for context, but are not explored in depth as they are beyond the scope of this review. Articles in which traits are observed sparsely over time, for example traits observed once at the beginning of an experiment and once after harvest, are excluded from this review, opting for consistent, relatively high temporal resolution tracking of traits over time.
Figure 2: Breakdown of publications reviewed, categorized by modeling approaches (left) and dataset types (right). The figure highlights the distribution of works across regression models, latent representation models, functional-structural plant models, and computer vision-based techniques, as well as their focus on temporal scalar, 2D image-based, and 3D point cloud datasets.
2.2 Plant Growth Representations and Datasets
@@
2.3 Limitations
This review highlight works that have contributed significantly towards advanced techniques in pattern recognition by providing deeper insights into the mathematical frameworks or, in the case of more traditional forecasting methods, are recent and representative of the fundamental principles with some novel insights discussed. Due to the exploratory nature of this review of novel and SOTA approaches within this developing field, the availability of works and datasets that cover longitudinal predictive forecasting and simulated growth representations and meet this criterion is limited. We have omitted some works that simply use models already established by other works, and thus the number of works available for a full systemic review is not possible. Moreover, this review is constrained to published research and datasets, which may lead to an inherent bias towards studies and datasets that have achieved publication rather than those that may be equally relevant but remain in preprint stages, unpublished, or private.
@@ 2.3 Limitations
Sun et al. (Soybean-MVS) [42] Soybean (Glycine max var. DN251, DN252, DN253, HN48, HN51) Segmented plant structure (leaf shape, stem, branches) Point cloud (3D)
Schunck et al. (Pheno4D) [43] Maize (Zea mays), Tomato (Solanum lycopersicum) Segmented plant structure (leaf shape, stem, branches) Point cloud (3D)
Sultan et al. (Shoots and Roots) [44] Maize (Zea mays), Tomato (Solanum lycopersicum), rice (Azucena, Bala), Arabidopsis (Arabidopsis thaliana) Root system traits and structure Point cloud (3D)
Uchiyama et al. (KOMATSUNA) [45] Komatsuna (Brassica rapa var. perviridis) Plant structure Multi-view RGBD images and annotated masks (2D/3D)
Jayasuriya et al. [46] Capsicum (Capsicum annuum L.) Plant structure RGBD images (2D/3D)
Kierdorf et al. (GrowliFlower) [45] Cauliflower (Brassica oleracea var. botrytis) Plant structure RGB and multi-spectral images (2D)
Shadrin et al. [47] Lettuce (Lactuca sativa) Leaf growth and leaf segmentation Top-down RGB images and annotated masks (2D), and environmental sensor readings (tabular)
Genze et al. [48] Maize (Zea mays L.), Sorghum (Sorghum bicolor) Plant structure Top-down RGB images and annotated masks (2D)
Thoday-Kennedy et al. [49] Safflower (Carthamus tinctorius L.) Plant morphology and salt response Multi-view RGB images (2D)
Lärm et al. [50] Maize (Zea mays), Wheat (Triticum aestivum) Root system traits and structure RGB images and ground-penetrating radar map (2D), soil sensor data (tabular)
Parmentier et al. [51] General vegetation in Arctic ecosystems Vegetative phenology and climate interactions RGB images (2D) and environmental vegetative index data (tabular)
Craig et al. [52] Maize (Zea mays L.) Various agronomic traits and vegetative index Scalar (tabular)
Dépigny et al. (Plantain-Optim) [53] Plantain (Musa acuminata × Musa balbisiana) Various agronomic traits and biomass Scalar (tabular)
Ngaba et al. [54] Sugarcane (Saccharum officinarum) Various agronomic traits and biomass Scalar (tabular)
Newman et al. [55] Multiple field-crop species Long-term agronomic traits and biomass Scalar (tabular)
Campbell et al. [56] Maize (Zea mays) Photosynthetic function and vegetative index Scalar (tabular)
Manjarrez-Sanchez et al. [57] Greenhouse Tomato (Lycopersicum esculentum) Photosynthetic function and agronomic traits Scalar (tabular)
Niu et al. [58] Maize (Zea mays) Long-term phenological and climate interactions Scalar (tabular)
Zhu et al. [59] Multiple field-crop species Long-term phenological and climate interactions Scalar (tabular)
Mergner et al. [60] Arabidopsis (Arabidopsis thaliana) Gene expression Scalar (tabular)
Castro et al. [61] Camu-camu (Myrciaria dubia) Gene expression Scalar (tabular)
Peng et al. [62] Peanut (Arachis hypogaea) Gene expression Scalar (tabular)